OpenStudy (shootforthestars):

put in slope intercept form 5x-4y=3 passes through (7,-5) put in slope intercept form 2x-5y=9 passes through (2,-9)

OpenStudy (*ar*):

r u trying to get y by itself ?

OpenStudy (*ar*):



OpenStudy (shootforthestars):

ya, but couldn't figure it out

OpenStudy (shootforthestars):

thats not it, because its passing through a point

OpenStudy (*ar*):

hold on i can do it

OpenStudy (*ar*):

r u trying to fine Find the Perpendicular Line

OpenStudy (shootforthestars):

no the slope intercept equation

OpenStudy (shootforthestars):

the answer is y=-4/5x - 19/5 but i do not know how to get to it

OpenStudy (zenmo):

what did you get for y when plugging in ( 7 , -5)?

OpenStudy (zenmo):

is that the answer to the first or 2nd problem?

OpenStudy (shootforthestars):

(-5)=-4/5(7) - 19/5 -5 = -28/5 - 19/5 -5 = -47/5 ?

OpenStudy (zenmo):

you need help on the first problem right?

OpenStudy (shootforthestars):

No, it was a problem I had to do in my class, but I got it wrong and I want to know how to solve it

OpenStudy (shootforthestars):

yup

OpenStudy (zenmo):

the answer to the first one is wrong, it is supposed to be \[y=\frac{ 5 }{ 4 }x-\frac{ 55 }{ 4 }\]

OpenStudy (zenmo):

\[5x-4y=3\] \[-4y=3-5x\] \[y=\frac{ 5 }{ 4 }x-\frac{ 3 }{ 4 }\]

OpenStudy (zenmo):

Looking at y=mx+b (slope intercept form), m = 5/4

OpenStudy (zenmo):

Plug m into the slope intercept form, \[y=\frac{ 5 }{ 4 }x+b\]

OpenStudy (zenmo):

\[-5=\frac{ 5 }{ 4 }(7) + b\] Plugging in ( 7, -5) in.

OpenStudy (shootforthestars):

-55/4?

OpenStudy (zenmo):

we need to isolate b, so subtract 35/4 to the other side of -5 to get \[-5-\frac{ 35 }{ 4 }=b\]

OpenStudy (zenmo):

yes and that is b.

OpenStudy (zenmo):

now, since we found b (b = -55/4) and the slope m ( m=5/4), plug it into the slope-intercept formula to get y=(5/4)x-55/4

OpenStudy (shootforthestars):

oh ok

OpenStudy (zenmo):

The purpose of the first part is to find the slope of m. Once you found the slope of m. Use the slope-intercept formula y=mx+b, plug in m, after plug in the given points to find b. Then after all is found, just plug m and b into y=mx+b.

OpenStudy (shootforthestars):

ok

OpenStudy (zenmo):

you said the answer was -4/5x-19/5?

OpenStudy (shootforthestars):

y=5/4x-55/4

OpenStudy (shootforthestars):

ya thats what my teacher said

OpenStudy (zenmo):

and that is for the first problem correct?

OpenStudy (shootforthestars):

yup

OpenStudy (zenmo):

I think your teacher made a mistake lol

OpenStudy (shootforthestars):

Well he marked it wrong on my test :(

OpenStudy (zenmo):

Show your work to another teacher, you are guaranteed to get your points back due to the teacher making a mistake.

OpenStudy (shootforthestars):

ok, thanks

OpenStudy (zenmo):

do you need help for the 2nd problem?

OpenStudy (shootforthestars):

No, I know how to do it now

OpenStudy (shootforthestars):

Thanks :D

OpenStudy (zenmo):

okay no problem, be sure to talk to that teacher or to some authority though. Those points taken off is invalid.